OpenStudy (anonymous):

Solve the following system of equations. 3x + 3y + 2z = –5 x + y + 3z = –4 2x + 2y + 5z = –7 I found z which = -1 but I keep somehow when trying to find x or y keep finding z and its ticking me off

OpenStudy (anonymous):

i dont understand what im doing wrong

OpenStudy (anonymous):

Yes z = -1

OpenStudy (anonymous):

ok so whats y or x... I keep coming up with the equation finding z. and Im trying to find y or x now

OpenStudy (anonymous):

Substitute z in the equations

OpenStudy (anonymous):

infinite solutions

OpenStudy (anonymous):

x and y take many values...

OpenStudy (anonymous):

satisfying x + y = -1 and z = -1

OpenStudy (anonymous):

so u have only two equations but 3 variables..in this case x and y take infinite no of values

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

@him1618 that's what I said

OpenStudy (anonymous):

its important to highlight the relation to the no of equations and variables

OpenStudy (anonymous):

the answers givin for this equation are a: (-4,-3,1) b:(-4,3,-1) c:(4,-3,-1) d:(-4,-3,-1)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

no infinite solutions ?

OpenStudy (anonymous):

x+y=-1

OpenStudy (anonymous):

As i said x and y should be satisfying x+y = -1 and z= -1....Only b satisfies these conditions

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

thank you , and yes!

OpenStudy (anonymous):

welcome:)